# SoK: A Literature and Engineering Review of Regular Expression Denial of Service (ReDoS)

### v3

@@ Appendix B ReDoS in Practice
This section provides a detailed account of the study described in Section 2.3.2. It outlines the methodology employed to investigate ReDoS vulnerabilities, including the criteria for selecting relevant data and the dimensions of analysis. Additionally, the section presents the key findings and results of the study
B.1. Methodology
B.1.1. ReDoS in the National Vulnerability Database
We performed a comprehensive analysis of NVD data from 2014 to 2023, focusing on CVEs associated with OWASP Top-10 and ReDoS-related vulnerabilities, using the classifications available on the OWASP website [^5] https://owasp.org/www-project-top-ten/. To ensure that ReDoS CVEs were accurately identified, we employed a systematic two-stage keyword search:
@@ B.1.1. ReDoS in the National Vulnerability Database
This method effectively identified mislabeled ReDoS CVEs, with our checks confirming no false positives in the sampled data. We further categorized the CVEs by their respective ecosystems, using specific keywords for each:
  • npm: npm, node.js, nodejs, javascript, js
  • PyPI: pypi, python, pip, django, flask
  • Gems: rubygems, ruby, gems, rails
  • Maven: maven, java, spring
  • Packagist: packagist, composer, php, laravel
  • Cargo: cargo, rust, crates.io, rust-lang
Our verification of a sample from each ecosystem showed no false positives, confirming the accuracy of our categorization.
B.2. Results
B.2.1. ReDoS in the National Vulnerability Database
Our analysis of CVE data from the NVD over the past 10 years has provided insights into the incidence of OWASP Top-10 and ReDoS-related CVEs. These trends are illustrated in Figure 6, with a specific focus on ReDoS-related CVEs presented separately in Figure 7. Additionally, we explored the distribution of OWASP Top-10 and ReDoS CWEs across different ecosystems, as shown in Figure 3. This visualization also includes the average occurrence of each CWE type throughout all ecosystems.
Figure 6. Number of OWASP Top-10 and ReDoS-related CVEs (2014-2023).
Figure 7. Number of ReDoS-related CVEs (2014-2023).
Appendix C wrk-redos Framework
The wrk-redos framework is designed to provide a practical, simulation-based approach to evaluate the severity of ReDoS vulnerabilities. Although many existing studies analyze ReDoS from a theoretical perspective, real-world scenarios often involve complex interactions between the regular expression engine, the application, and the underlying infrastructure. wrk-redos addresses this gap by offering tools and methodologies to simulate realistic attack scenarios, allowing developers and researchers to understand the impact of ReDoS vulnerabilities on system performance.
This framework is beneficial for the following purposes:
  • Benchmarking quality of service (QoS) degradation during a ReDoS attack.
  • Validating the severity of vulnerabilities found in regexes, especially in scenarios with third-party dependencies.
  • Providing actionable insights to mitigate the risks associated with vulnerable regex patterns.
The following subsections outline the use cases, threat model, and implementation details of the framework (Appendix C.1), along with the results of simulations (Appendix C.2) and an analysis of their applicability in realistic environments (Appendix C.3).
C.1. Realistic Simulation of ReDoS Attacks
As shown in Section 3, academic work on ReDoS can benefit from more realistic threat models and availability-oriented evaluations, to better communicate the ReDoS-associated risks. To this end, we present a framework for assessing the severity of vulnerable regular expressions via attack simulations. We first define the main considered use cases, the threat model and the ReDoS definition, then present wrk-redos, our prototype for performing measurements in the proposed framework.
Use Cases. As discussed in the introduction, developers are often overwhelmed by the number of supposedly-vulnerable regular expressions in their code base, in particular, in code they did not write, e.g., shipped with third-party dependencies. To confirm the severity of such security findings, we propose simulating actual ReDoS attacks and benchmark the corresponding degradation in quality of service, as perceived by concurrent users. In this way, developers can understand the attacker’s leverage in an asymmetric DoS, enabled by the problematic regular expression.
Considering the current state of regex engines discussed in Section 4, we note a second usage scenario for our framework. It can assist developers with understanding the effect of migrating their application to a different engine version, from a ReDoS resilience perspective. Concretely, developers can perform attack simulations before and after changing the runtime version and measure how the attacker’s advantage changes as a result of this migration. Since regex engine migration carries risks, notably that of semantic changes, wrk-redos can help engineers assess the risk-reward tradeoff here.
Threat Model. We assume that attackers and users can interact simultaneously with the target system via two endpoints: one that is benign and one that is vulnerable to ReDoS attacks. Attackers have limited bandwidth and know how to create ReDoS payloads, but we do not assume any special deployment features like redundancy or intrusion detection systems. This model restricts attackers from sending excessively large payloads and does not presume a specific regex engine type, which aligns more closely with realistic scenarios. Under this threat model, we propose an alternative definition for ReDoS vulnerabilities that we believe is more in line with the practitioners’ expectations:
Definition 0.
A problematic regular expression is considered a vulnerability if and only it is accompanied by an attack simulation with the following characteristics:
  • C1: The vulnerability is exercised in a realistically-deployed web application, e.g., in the cloud.
  • C2: The attackers’ bandwidth is restricted to a reasonably low value, in line with the asymmetric setup.
  • C3: The simulation shows that concurrent benign users experience dropped requests or extremely high latency.
This definition has certain advantages over previous ones proposed in the literature. It discards minor slowdowns caused by ReDoS issues, which cannot be leveraged for an asymmetric attack. It emphasizes the server’s QoS degradation, as perceived by benign users, instead of merely measuring the slowdown produced by individual payloads. Moreover, it requires critical availability consequences like dropped requests for an issue to be considered a vulnerability. We believe that this will help alleviate the security fatigue regarding ReDoS and prioritize harmful ReDoS issues. Most importantly, our definition mandates that the web application needs to be deployed in a setup that closely resembles a real production environment. In this way, certain deployment-time components such as the reversed proxy might absorb some of the malicious ReDoS traffic and inadvertently prevent the attack.
Figure 8. Overview of wrk-redos, an attack simulation framework for a realistic assessment of ReDoS findings.
Simulation Framework. Considering the criteria we established for a successful attack, we propose wrk-redos a simulation tool that measures the effectiveness of ReDoS attacks leveraging a target regular expression. Figure 8 depicts how wrk-redos operates. wrk-redos takes the user’s simulation configuration, such as attack bandwidth and window, measurement bandwidth, and experiment duration. It then launches two concurrent workloads: (i) a benign one that is triggered first and continuously benchmarks the target server, i.e., sends high-bandwidth measurement requests, (ii) a malicious one that sends ReDoS payloads in a restricted time window, at a predefined rate. wrk-redos communicates with these processes, manages the data configuration, and gathers output metrics, such as throughput, latency, and errors for benign users. Adopting a remote benchmarking approach, wrk-redos highlights the attack’s impact on benign users (C3), measuring the system’s capability to handle both legitimate and malicious traffic concurrently. Nonetheless, to comply with the other two simulation characteristics in the definition above, users of wrk-redos must ensure that they configure attacks with sufficiently low bandwidth (C2) and that the target application is deployed under realistic conditions (C1). Similar to related work such as Rampart (Meng et al., 2018), Shan et al. (Shan et al., 2017), and Olivo et al. (Olivo et al., 2015a), our framework performs short attack bursts and observes the server’s behavior before and after the attack. This method enables asymmetric attacks, similar to low-rate denial of service and Shrew attacks(Zhijun et al., 2020), both known to bypass existing DoS defenses. Moreover, wrk-redos utilizes wrk2 (ten, 2024), a popular HTTP benchmarking tool (Fuerst et al., 2020; Unnikrishnan, 2023; Ding et al., 2023; Gouveia et al., 2020; Jacquet et al., 2023; Maissen et al., 2020) to simulate both benign and malicious traffic. This allows us to maximize the amount of traffic sent from one machine and reuse a mature measurement infrastructure adopted both by practitioners and academics. wrk2 follows standards set in benchmarking literature to ensure accurate, stable, and relevant measurements.
C.2. Experiments
Table 7. Performance metrics for 10 NPM and PyPI libraries at input sizes 25, 1000, and 5000. Metrics include 99.99th percentile latency ($p_{99.99}$), mean ($\bar{x}$), timeout requests ($N_{u--sw}$), and failed requests ($N_{failed}$).
Library Input size = 25 Input size = 1000 Input size = 5000 Input size = 5000 (RE2) $p_{99.99}$ $\bar{x}$ $N_{u--sw}$ $N_{failed}$ $p_{99.99}$ $\bar{x}$ $N_{u--sw}$ $N_{failed}$ $p_{99.99}$ $\bar{x}$ $N_{u--sw}$ $N_{failed}$ $N_{u--sw}$ $N_{failed}$ black 561.49 128.66 0 0 178.27 132.70 3401 (75.5%) 50 777.65 136.52 3251 (72.2%) 50 247 (13.7%) 0 celo 173.76 124.52 3401 (75.5%) 50 412.93 111.14 3401 (75.5%) 50 151.77 128.00 3301 (73.3%) 50 0 0 django-truncate 438.82 113.41 0 0 479.60 154.64 0 0 12801.15 2615.76 0 0 N/A N/A django-intcomma 426.59 111.62 0 0 189.72 122.68 0 0 7332.25 1121.46 0 0 N/A N/A semgrep 9644.78 1675.03 0 0 459.06 120.57 3301 (73.3%) 50 543.81 139.27 3396 (75.4%) 50 0 0 charset 574.81 129.82 0 0 483.45 129.83 0 0 5496.47 134.67 0 0 0 0 node-email-check 342.45 137.12 3301 (73.3%) 50 487.60 133.98 0 0 621.69 117.48 0 0 0 0 cookiejar 704.84 122.32 0 0 185.87 118.65 0 0 5238.85 847.67 0 0 0 0 lodash 437.11 112.79 0 0 539.86 129.12 0 0 197.13 126.80 0 0 0 0 semver 652.86 110.85 0 0 1407.34 164.78 0 0 10154.02 1797.97 0 0 0 0
Table 8. Overview of npm and PyPI libraries used in the study, including their versions, associated CVEs, programming languages, and severity scores.
Library Version CVE Language Severity
black 24.1.0 CVE-2024-21503 Python 5.3
celo 1.0.0a5 CVE-2022-42966 Python 5.9
django-truncate 3.2.0 CVE-2023-43665 Python 5.9
django-intcomma 3.2.0 CVE-2024-24680 Python 5.9
semgrep 1.23.0 CVE-2023-32758 Python 7.5
charset 1.0.0 CVE-2017-16098 Nodejs 7.5
node-email-check 1.0.0 CVE-2023-39619 Nodejs 7.5
cookiejar 2.1.3 CVE-2022-25901 Nodejs 5.3
lodash 4.17.16 CVE-2020-28500 Nodejs 5.3
semver 5.1.1 CVE-2022-25883 Nodejs 5.3
To illustrate the utility of wrk-redos, we analyzed previously reported ReDoS vulnerabilities in open-source libraries by conducting equivalent attacks of varying intensities. We selected ten libraries from the npm and PyPI ecosystems, all with reviewed vulnerability reports in the GitHub Advisory Database. These libraries were chosen based on their severity scores and available proof-of-concept payloads. Details about the tested libraries and their versions are provided in Table 8. Despite their popularity—eight libraries with over 850,000 weekly downloads and seven with more than 3.5 million—these libraries are still susceptible to security issues. We used both Heroku’s basic dyno (512 MB memory, 1 CPU) and an AWS EC2 instance (1 GB memory, 1 CPU) to set up our sample applications. The EC2 instance ran custom Docker applications to implement the second use case described in Section C.1, which involved replacing Python’s re module with RE2 and Node.js’s regex engine with RE2, instead of the default V8 engine. However, we could not perform the same measurements for django-intcomma due to a stack overflow caused by RE2’s internals. This setup enabled us to simulate and measure the impact of migrating to a ReDoS-safe engine, comparing ReDoS resilience and performance before and after the engine swap. We created Express and Flask applications to process user input through these vulnerable libraries, with dedicated API endpoints for each. Malicious payloads were sent via headers to test the servers’ vulnerability to attacks.
Figure 9. Throughput and latency for benign users for django-truncate library at payload sizes 25, 1000, 5000, and 10000.
Table 7 presents the results of our simulation experiments using three payload sizes (25, 1000, and 5000 characters), with an attack burst lasting 10 seconds at a rate of 100 req/s, and a measurement window of 60 seconds at the same rate. We calculate and list various metrics that demonstrate the resilience of the target systems under these attacks. As the intensity of the attack increases, we observe a consistent rise in latency metrics (average and tail). Additionally, the last two columns of the table show the number of timeouts and failed requests with the 5000-character payload when RE2 was enabled in both Node.js—V8 and Python. Despite charset having a high severity score of 7.5, its impact on performance metrics such as average and median latencies was minimal, with increases only in the millisecond range except for some outliers at the $99.99th$ percentile ($p_{99.99}$). Conversely, semver, with a lower severity score of 5.3, seriously reduced server throughput to less than 10% with an input size of 5000 characters and significantly increased the average latency of benign requests for 1000 characters. These observations suggest that the severity scores provided in advisory databases might not consistently reflect their effect on a system’s vulnerability to DoS attacks. Additionally, different advisory databases may assign varying severity scores to the same vulnerability, as seen with node-email-check, where GitHub Advisory Database rates it as a high vulnerability (7.5), whereas Snyk.io considers it medium severity (5.3). Interestingly, celo exhibited a complete server shutdown with an input length of only 25 characters but showed no adverse effects with larger inputs of 1000 and 5000 characters. This resilience at larger sizes can be attributed to the input restriction implemented in the celo library, which sanitizes input by ensuring the email length is greater than 5 and less than 61 characters, as shown in Figure 10. While this approach effectively filters out malicious inputs longer than 61 characters and prevents large-payload attacks, it does not prevent Regular Expression Denial of Service (ReDoS) attacks with smaller, targeted inputs exploiting the vulnerable expression.
1function validateEmail (email) {
2 if (typeof email === ’string’ && email.length > 5 && email.length < 61 && validateRegex.test(email)) {
3 return email.toLowerCase();
4 } else {
5 return false;
6 }
7}
Figure 10. Email validation code snippet from Celo library
Let us now consider the case of django-truncate in more detail. Figure 9 illustrates the throughput and latency of the Django application under various attack scenarios. In Django version 3.2, the method django.utils.text.Truncator is vulnerable to ReDoS, when used with html=True. In our setup, malicious data is sent by including it in an HTTP header — a common method for data transmission to the server — to exploit this vulnerability. The graph depicts the server’s response to different simulated attack scenarios with payloads of lengths 25, 1000, 5000, and 10000. We observe an increase in both average and median latency as the payload sizes increase from 25 to 1000, indicating a degradation in server performance during the attack. Moreover, with a payload of 5000, there is a significant reduction in server throughput, approximately to 10%, which signifies a substantial impact of the attack. However, this throughput rebounds to baseline levels shortly after the cessation of the attack. Conversely, the 10000 payload scenario deviates from the expected outcome, exhibiting no discernible effect on server performance. This anomaly arises because the payload size exceeds Heroku’s header limit, resulting in a rejection of the malicious requests by the server’s infrastructure. Thus, despite the attacker dispatching larger payloads, the server remains unaffected, due to an intrinsic safeguard against excessively large payloads within the platform. This supports our claim that ReDoS findings should be evaluated in realistic scenarios, where such complexities are considered.
C.3. Analysis of Simulation Results
When simulating attacks like ReDoS, it is essential to employ a multi-faceted measurement strategy that goes beyond single metrics such as tail latency. This avoids the over-simplification trap of a single viewpoint, as discussed below.
Detailed Analysis of Metrics. Let us consider the case of semgrep in Table 7 to illustrate how relying solely on tail latency can sometimes paint an incomplete picture of an attack’s true impact. With input sizes of 25 versus 1000, the data suggests a disproportionate effect on server performance at smaller input sizes. Initially, it appears more disruptive due to higher latency spikes. Yet, this view does not investigate the server’s ability to recover quickly once the attack ceases, which is crucial for assessing the system’s resilience.
As shown in Figure 11, the effect of a ReDoS attack on the semgrep library using input sizes of 25 and 1000 characters highlights this issue. For an input size of 25, the throughput decreases to almost 10%, but it recovers immediately once the attack ceases. However, with an input size of 1000, the server fails to recover within the experiment window, demonstrating a significant vulnerability to larger ReDoS attacks. Even though the reported tail latency is lower due to incomplete data, the actual impact is significantly more severe since the server remains non-operational.
Figure 11. Server throughput over time for input sizes of 25 and 1000 using the Semgrep library.
Figure 12. Throughput and latency for benign users at 500 req/s bandwidth. The increasing latency and fluctuating throughput indicate bandwidth exceeding server capacity, leading to an inadvertent DDoS attack.
Effects of Payload Size. We note that there is often a key threshold beyond which increasing the payload size has no additional negative impact. For instance, comparing input sizes of 1000 and 5000 for semgrep in Table 7 shows negligible differences in the observed metrics. This indicates that once a payload is sufficiently large to overwhelm the server, further increases do not exacerbate the attack’s effect.
Setting Appropriate Bandwidth. When evaluating the severity of a ReDoS exploit, it is crucial to ensure that the simulation itself does not inadvertently launch a traditional DoS attack on the server. This can be avoided by properly setting the attack bandwidth in the benchmarking tool; otherwise, the results fail to show the effect of the ReDoS payload, but of the high-bandwidth traffic. Figure 12 illustrates this issue, showing the consequences of running wrk-redos at an attack rate of 500 requests per second, which exceeds the server’s capacity. This is indicated by the fluctuations in throughput and the increase in latency even before the actual attack begins. As a result, the observed performance degradation can be attributed mainly to the high traffic volume, instead of the ReDoS payloads. Therefore, it is essential to calibrate both the benign and attack bandwidths to accurately assess the impact of a ReDoS attack.
Stable Measurement Conditions. Preparing and calibrating the system before conducting attack simulations is essential. This involves stabilizing both throughput and latency with benign loads to ensure that the baseline performance is reliable and reflective of typical operational conditions. wrk-redos facilitates this by simulating benign conditions, at first, to verify the system’s stability before launching the attack simulation, helping to ensure that the subsequent measurements accurately reflect the attack’s impact.
Appendix D Other Tables and Figures
Table 9. Regex engine implementations selected for evaluating ReDoS defenses
JavaScript Node.js—V8 (Nod, 2024; V8V, 2024) Deno, Bun, SpiderMonkey, JavaScriptCore Dominates server-side JavaScript usage (98.8% market share) (Dis, [n. d.])
Ruby MRI/CRuby (Rub, 2024) JRuby, Rubinius, mruby, RubyMotion Reference implementation (Rub, [n. d.])
C# .NET (Dot, 2024) Mono Reference implementation (Lan, 2024)
Java OpenJDK (Ope, 2024) Amazon Corretto, Azul Zulu, Liberica JDK, Eclipse Adoptium… Most widely used JDK of 2024 with a 20.8% usage rate (202, [n. d.])
PHP Zend Engine (Php, 2024) HHVM, PeachPie, Quercus, Parrot Standard PHP interpreter is driven by the Zend Engine (PHP, [n. d.]; php, [n. d.])
Perl perl5 (Per, 2024) N/A (for Perl 5) Only implementation for Perl 5 (Per, [n. d.])
Rust rustc (Rus, 2024b) mrustc, gccrs Official and only fully functional compiler for Rust (rus, [n. d.])
Go gc (Gol, 2024) gccgo, gofrontend, TinyGo, GopherJS, yaegi Official compiler included in Go releases (Int, [n. d.])
Python CPython (Pyt, 2024) PyPy, Stackless Python, MicroPython, CircuitPython, IronPython, Jython Reference implementation (Pyt, [n. d.])
PL: Programming Lanuage, Impl.: Implementation (runtime environment, engine, interpreter, or compiler)
Table 10. List of all works included in this SoK, sorted by year of publication.
Paper Year Venue Kirrage et al. (Kirrage et al., 2013) 2013 NSS Weideman et al. (Weideman et al., 2016) 2016 CIAA Davis et al. (Davis et al., 2017) 2017 EuroSec Rathnayake et al. (Rathnayake and Thielecke, 2017) 2017 arXiv Wüstholz et al. (Wüstholz et al., 2017) 2017 TACAS Petsios et al. (Petsios et al., 2017) 2017 CCS Staicu et al. (Staicu and Pradel, 2018) 2018 USENIX Security Davis et al. (Davis et al., 2018a) 2018 ESEC/FSE Shen et al. (Shen et al., 2018) 2018 ASE Noller et al. (Noller et al., 2018) 2018 ISSTA Davis et al. (Davis et al., 2018b) 2018 USENIX Security Davis et al. (Davis et al., 2019) 2019 ESEC/FSE Demoulin et al. (Demoulin et al., 2019) 2019 USENIX ATC Li et al. (Li et al., 2020) 2020 ASE Turoňová et al. (Turoňová et al., 2020) 2020 OOPSLA Liu et al. (Liu et al., 2021) 2021 IEEE S&P Li et al. (Li et al., 2021) 2021 USENIX Security Bai et al. (Bai et al., 2021) 2021 IEEE S&P Davis et al. (Davis et al., 2021) 2021 IEEE S&P Turoňová et al. (Turoňová et al., 2022) 2022 USENIX Security Barlas et al. (Barlas et al., 2022) 2022 ICSE Kluban et al. (Kluban et al., 2022) 2022 AsiaCCS McLaughlin et al. (McLaughlin et al., 2022) 2022 USENIX Security Chida et al. (Chida and Terauchi, 2022) 2022 IEEE S&P Li et al. (Li et al., 2022) 2022 USENIX Security Bhuiyan et al. (Bhuiyan et al., 2023) 2023 ICSE Wang et al. (Wang et al., 2023) 2023 IEEE S&P Parolini et al. (Parolini and Miné, 2023) 2023 Science of Computer Prog. Hassan et al. (Hassan et al., 2023) 2023 IEEE S&P Tandon et al. (Tandon et al., 2023) 2023 RAID Su et al. (Su et al., 2024) 2024 USENIX Security Siddiq et al. (Siddiq et al., 2024a) 2024 ICPC Siddiq et al. (Siddiq et al., 2024b) 2024 ICSE Kluban et al. (Kluban et al., 2024) 2024 TOPS
D.1. Regex Engine Selection Summary
Table 9 provides a comprehensive overview of programming language, hence the regex engine implementations selected for our ReDoS defenses review (Section 4). For each programming language considered in our study, we list the chosen implementation, its alternatives, and the rationale behind our selection. In summary, this table reflects our effort to analyze the most relevant and widely-used regex engines in production environments, focusing on reference implementations or those with significant usage rate compared to alternatives where possible.
D.2. Included Works in the Systematization
Table 10 presents all the works included in the literature review in Section 3 of this SoK. The papers are sorted by year of publication and also include the publishing venue or conference. These works were selected based on the inclusion criteria described in Section 3, ensuring relevance to ReDoS-specific research. The selected papers represent the state-of-the-art in the field, covering key aspects such as detection, prevention, and mitigation strategies, and reflecting significant contributions to advancing the understanding of ReDoS.
D.3. Excluded Works in the Literature Review
Some closely related papers were excluded from the literature review because they do not focus on ReDoS-specific vulnerabilities or their evaluation lacks any ReDoS-specific aspects. These exclusions were necessary to ensure that the review remains precise and relevant to the study of ReDoS. For instance, (Cox, 2007; Varatalu et al., 2025; Turoňová et al., 2022; Shan et al., 2017) were excluded because they focused on broader denial-of-service (DoS) vulnerabilities or algorithmic complexity attacks. Other papers were excluded for evaluating performance issues in regex engines without demonstrating their impact on ReDoS vulnerabilities.
@@ D.3. Excluded Works in the Literature Review
Table 11. Example works excluded with explanation
Paper Reason for Exclusion
Cox, R. (Cox, 2007) Evaluation does not address ReDoS-specific scenarios.
Turoňová et al. (Turoňová et al., 2022) Evaluation does not address ReDoS-specific scenarios.
Shan et al. (Shan et al., 2017) Addresses general denial-of-service (DoS) attacks.
Wei et al. (Meng et al., 2018) Addresses general denial-of-service (DoS) attacks.
D.4. Analyzing Developer Discussions on GitHub ReDoS Issues
To analyze discussions around ReDoS vulnerabilities, we utilized the GitHub REST API to search for issues containing the term redos in their title, body, or comments. This approach enabled the identification of both open and closed issues across various repositories, capturing a wide range of cases, including those actively being discussed and those already resolved. To focus on meaningful engagement, we filtered the results to include only issues with at least two comments, ensuring the dataset represented discussions with substantive input. After retrieval, we manually reviewed the comments within these issues to identify and extract relevant discussions about ReDoS vulnerabilities.
Table 12 showcases 10 examples of comments from selected issues, providing URLs for further exploration. These discussions often reveal that developers do not perceive ReDoS vulnerabilities as relevant or critical, leading to delays or reluctance in addressing the issues. While some developers prioritize security, others downplay the severity of these vulnerabilities, resulting in limited attention or resolution efforts. These trends underscore the challenges faced in addressing ReDoS issues in open-source projects.
Table 12. Selected ReDoS-Related Comments and URLs

### v4

@@ Appendix B ReDoS in Practice
We conducted a comprehensive analysis of NVD CVE data from 2014 to 2023 in order to quantify ReDoS vulnerabilities in practice. As part of this effort, we utilized the CWE field within the CVE reports. For each OWASP Top 10 category, we counted the number of CVEs that included a CWE entry associated with that specific OWASP category.
To identify ReDoS CVEs, we employed a systematic two-stage keyword search:
@@
This method effectively identified ReDoS-related CVEs, and our verification of a sample confirmed no false positives.
Appendix C Works Excluded From the Literature Review
Some closely related papers were excluded from the literature review because they do not focus on ReDoS-specific vulnerabilities or their evaluation lacks any ReDoS-specific aspects. These exclusions were necessary to ensure that the review remains precise and relevant to the study of ReDoS. For instance, (Cox, 2007; Varatalu et al., 2025; Shan et al., 2017) were excluded because they focused on broader DoS vulnerabilities or algorithmic complexity attacks. Other papers were excluded for evaluating performance issues in regex engines without demonstrating their impact on ReDoS vulnerabilities.
@@ Appendix C Works Excluded From the Literature Review
Appendix D Additional Details of the Engineering Review
This appendix provides supplementary information to support the engineering review presented in §4. It includes details on the regex engine selection process and the engine versions used in our measurements.
D.1. Regex Engine Selection
Table 8 summarizes the regex engine implementations reviewed in our ReDoS defenses review (§4). For each programming language considered in our study, we list the selected implementation, notable alternatives, and our rationale. We prioritized (1) engines widely used in production settings (e.g., Node.js—V8, used by 98.8% of websites employing server-side JavaScript (Q-Success, 2024)); and (2) reference implementations with publicly accessible source code. These are typically cited in official language documentation and serve as the standard for the language (e.g., CPython for Python). These choices ensure both real-world relevance and access to implementation details for assessing ReDoS defenses.
D.2. Versions Used in the Measurements
To ensure the reproducibility of the experiments described in §4.2, Table 9 indicates the specific versions of the programming language runtimes and their regex engines that were evaluated. The old versions correspond to earlier releases that predate known or likely ReDoS-relevant mitigations, while the new versions refer to more recent releases that incorporate possible ReDoS defenses or performance improvements in their regex engines.
Appendix E Analyzing Developer Discussions on GitHub ReDoS Issues
In §5, we discussed the idea of “Developers push back”. We referred to conversations on GitHub and sites of other engineering discourse.
To find discussions about ReDoS vulnerabilities, we used the GitHub REST API to search for issues containing the term redos in their title, body, or comments. This approach helped us find both open and closed issues across many repositories, covering a range of cases, including those still active and those already resolved. We started with $\sim$46.4K issues and filtered them to include only those with at least two comments, leaving $\sim$7.4K issues. This helped us focus on discussions with more input. Out of these, $\sim$22.6K issues were not opened by bots. $\sim$4K of those are open, and $\sim$18.6K are closed. We manually reviewed 250 recent issues from each set, along with their comments, to find relevant ReDoS discussions.
Table 10 showcases 10 examples of comments from selected issues, providing URLs for further exploration. These discussions often reveal that developers do not perceive ReDoS vulnerabilities as relevant or critical, leading to delays or reluctance in addressing the issues. While some developers prioritize security, others downplay the severity of these vulnerabilities, resulting in limited attention or resolution efforts. However, we acknowledge this aspect of our investigation is incomplete and bears further study.
Table 10. Selected comments regarding ReDoS and corresponding issue URLs.